OpenStudy (anonymous):

How much would $300 invested at 4% interest compounded monthly be worth after 8 years? Round your answer to the nearest cent. here's the equation: A(t)=P(1+r/n)^nt i just don't know what to put where please help

OpenStudy (anonymous):

@Hero @mathstudent55

hero (hero):

What I mean to ask is, what do you think each variable in the formula means?

hero (hero):

Obviously t is in years, so 8 = t. You probably know that much at least.

OpenStudy (anonymous):

is it a(8)=300(1+.06/1)^8 ??

OpenStudy (anonymous):

@Hero

hero (hero):

The 300 and the 8 is correct.

OpenStudy (anonymous):

is that is?

OpenStudy (anonymous):

it?*

hero (hero):

There's more to it of course.

hero (hero):

Okay, so the P is for the principal, which you figured out is $300. The r is the interest rate which is 4% or .04. And the n represents number of times compounded. You figure that out by multiply 8 years times 12 months per year. So n = 96.

OpenStudy (anonymous):

so 8=300(1+.04/1)^96?

hero (hero):

Where do you get the "1" underneath the fraction? Isn't that supposed to be "n". What did we say n equals?

hero (hero):

Actually, n = 12, not 96

OpenStudy (anonymous):

oh oops. 8=300(1+.04/12)^96 ?

hero (hero):

So n represents number of times compounded per year which is 12. So A(8) = 300(1.04/12)^{96} is correct.

OpenStudy (anonymous):

is it 1.04 for the numerator? or 1+ .04/12?

hero (hero):

Actually you have it right it's 1 + .04/12

hero (hero):

I admit, I'm a little distracted atm. Not fully paying attention.

OpenStudy (anonymous):

okay. how do i solve it?

hero (hero):

At this point, it's just plug and play if you have the right calculator you should be able to calculate A(8).

OpenStudy (anonymous):

am i solving for a?

hero (hero):

You're finding A(8). Everything on the right side of the equation only has to be calculated.

OpenStudy (anonymous):

\[A(8)=300(1+\frac{ .04 }{ 12 })^96\]

OpenStudy (anonymous):

like this? i tried to put the 6 up next to the 9 but it didn't work

OpenStudy (anonymous):

@Hero

hero (hero):

Use brackets

hero (hero):

\(A(8)=300\left(1+\dfrac{ .04 }{ 12 }\right)^{96}\)

OpenStudy (anonymous):

i keep getting 51.61

OpenStudy (anonymous):

the answer choices are a) $412.92 b) $410.57 c) $308.10 d) $484.25

OpenStudy (anonymous):

do you get 51.61?

OpenStudy (anonymous):

@Hero

hero (hero):

Hang on

hero (hero):

I get 412.918...

OpenStudy (anonymous):

oh! okay. i see how you got that. thank you!!

hero (hero):

You're Welcome bro.